OpenStudy (anonymous):

Find all real solutions of the equation x^2-125=0. x_1 = ___ and x_2 = ____ with x_1

OpenStudy (anonymous):

First add 125 to both sides then x^2 = 125. Then use a radical on both sides which gives you 5 radical 5. This can be plus or minus 5radical5

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[x^{2}-\left( 5\sqrt{5} \right)^{2}=0,\left( x+5\sqrt{5} \right)\left( x-5\sqrt{5} \right)=0\] from the two find x